Real life example of three fundamental laws of chemistry?
1
Expert's answer
2020-10-19T14:05:59-0400

1.

Law of conservation of mass


The law of conservation of mass states that the total mass present before a chemical reaction is the same as the total mass present after the chemical reaction.


Example:

Let 12 g of carbon is allowed to burn by 32 g of oxygen. The product CO2 produce has the mass 44g which is the sum of mass of carbon and oxygen involved in the reaction.

I.e 44 g = (12+32) g


2.

Law of Definite Proportions


Law of Definite Proportions states that in a given type of chemical substance, the elements are always combined in the same proportions by mass.


Example:

If 2 gram of Hydrogen reacts with 16 grams of Oxygen, then by the Law of Definite Proportions, 4(=2×2) grams of Hydrogen must react with 32 (=2×32) grams of oxygen.


3.

The law of multiple proportions


The law of multiple proportions states that if two elements form more than one compound between them, the masses of one element combined with a fixed mass of the second element form in ratios of small integers.


Let's for every 1 g of compound A there are 0.571 g of oxygen and 0.429 g of carbon. The mass of oxygen per gram carbon is,

0.571g oxygen/0.429g carbon=1.33g oxygeng carbon


Similarly, for 1 g of compound B, there are 0.727 g oxygen and 0.273 g of carbon. The ration of mass of oxygen per gram of carbon is,

0.727g oxygen/0.273g carbon = 2.66g oxygeng carbon


Dividing the mass of oxygen per g of carbon of the compound B.

2.66/1.33=2


Hence ,the masses of oxygen combine with carbon in a 2:1 ratio which s consistent with the Law of Multiple Proportions since they are whole numbers.